OpenStudy (aravindg):

31. Show that each one of the following is an odd function: (i) f(x) = x3 - x for all x Є R (ii) f(x) x = 1/x for all x Є R – {0} (iii) f(x) = І x І/x for all x Є R – {0} (iv) f(x) = x3 + 3 x – 4 sin 3 x for all x Є R.

OpenStudy (anonymous):

same way if f(x)=f(-x) then it is even hence not odd

OpenStudy (zarkon):

show f(-x)=-f(x)

OpenStudy (zarkon):

functions don't have to be odd or even. They can be neither.

OpenStudy (zarkon):

so for the first one \[f(-x)=(-x)^3-(-x)=-x^3+x=-(x^3-x)=-f(x)\]